# Magic tricycles: efficient magic state generation with finite block-length quantum LDPC codes

### v1

@@ Appendix A Code construction and properties
This homological viewpoint is especially useful for constructing transversal logical gates. In the next section, we apply the formalism of [16] to equip the complex Eq.˜18 with an algebraic structure that enables transversal $CCZ$ circuits.
Appendix B Cup products and transversal CCZ action
First, we introduce the necessary formalism at a high level. A cup-product is an algebraic structure that can be defined on some cochain complexes, such as the one in Eq.˜18. Let us label the cochain spaces of the complex in Eq.˜18 as $0,1,2,3$-cochains from left to right, indicated with the notation $C^{0}$, $C^{1}$, $C^{2}$, and $C^{3}$ respectively. Furthermore, we are interested in the case where the $C^{i}$ are finite-dimensional vector spaces over $\mathbb{F}_{2}$ with bases $X^{i}$. We define the weight $|c|$ of a cochain $c\in C^{i}$ as the number of non-zero terms in its basis expansion over $X^{i}$. For example, for the tricycle code complex in Eq.˜18, the bases $X^{i}$ are defined by $G$, and the cochain weight coincides with the weight of group-algebra elements. We will always associate qubits with the $1$-cochains of a complex, with $0$- and 22- cochains corresponding to spaces of $X$ and $Z$ checks respectively. Thus, we will label the elements of the basis of 11-cochains $X^{1}$ by qubit labels $q_{1}\cdots q_{N}$ for the $N$ qubits of a CSS quantum code.
A cup-product is then an associative, bilinear map
$\cup:C^{i}\times C^{j}\rightarrow C^{i+j}$ (19)
with the convention that $C^{i+j}=0$ if $i+j>3$ in the case of the tricycle complex. Not every such associative bilinear map defines a cup-product; the map $\cup$ must satisfy an additional technical condition known as a Leibniz rule – see [41, 16, 100]. However, given such a cup-product, we can construct a circuit of diagonal physical gates that preserves the code-space of the CSS code associated with the cochain complex. In general, a cup-product on an $m$-term cochain complex will yield a physical circuit of diagonal $C^{m-2}Z$ gates from the m1m-1-th level of the Clifford hierarchy [16]. In the situation of interest, we have the 44-term cochain complex in Eq.˜18, leading to a circuit of $CCZ$ gates. As discussed in Sec.˜2.2, such a circuit can be specified by a trilinear function $f$ acting on the space of physical qubits – the $1$-cochains $C^{1}$ in our convention. The circuit associated with the cup-product is then defined by the function
$f_{\cup}(q_{i},q_{j},q_{k})=|q_{i}\cup q_{j}\cup q_{k}|\in\mathbb{F}_{2}$ (20)
where non-zero values indicate a physical $CCZ$ gate acting between the qubits of three distinct code-blocks labeled by those arguments of $f_{\cup}$. Circuits defined by such functions constructed from cup-products of finitely many 11-cochains (three, in this case) lead to sparse, finite depth circuits [16, 55]. Moreover, such functions $f_{\cup}$ naturally have the property that they vanish on 11-coboundaries: $f_{\cup}$ vanishes if any of its three arguments is of the form $d^{0}(c)$ for $c\in C^{0}$, where $d^{0}$ is the coboundary map from $C^{0}$ to $C^{1}$ – equivalent to the matrix $H_{X}^{T}$. This condition is thus equivalent to $f_{\cup}$ vanishing when any of its arguments corresponds to an $X$-type stabilizer (see Sec.˜2.2) and is called coboundary invariance. The coboundary invariance property is essential to ensure that the circuit defined by $f_{\cup}$ descends to a well-defined logical operation [55, 16] – intuitively, this is so that the action of $f_{CCZ}$ on logical operators does not change upon adding stabilizers to the logical operators.
The formalism of [16] constructs cup-products on classical $2$-term cochain complexes which satisfy certain sufficient conditions such that homological and balanced product complexes constructed from them (see supplementary material) are also equipped with an appropriate cup-product. This yields a trilinear function of the form in Eq.˜20 on the product complexes associated with quantum CSS codes. Since tricycle codes are balanced products of classical Abelian group-algebra codes, these methods readily apply, which we review here. Note that for this to work, we require a mild technical condition that the check-weight of the classical codes is even (see [16] for details) – this is why we do not consider tricycle codes that are balanced products of classical group-algebra codes with weight $3$ checks despite the lower resulting check weight.
To equip a two-term cochain complex with a cup-product, one must construct partitions of the 11-coboundaries known as pre-orientations. For a classical Abelian group-algebra code associated with the two-term cochain complex of Eq.˜16 defined by an element $a\in\mathbb{F}_{2}[G]$, this amounts to choosing a partition of the group algebra element into smaller disjoint group-algebra elements
$a=a_{in}+a_{out}+a_{free}$ (21)
where $a_{in},a_{out},a_{free}\in\mathbb{F}_{2}[G]$. Different choices of partitions induce different cup-products on the classical code according to the rules in definition 5.2 of [16]. Given three such classical group algebra codes defined by elements $a,b,c\in\mathbb{F}_{2}[G]$, each with their respective in, out, and free partitions, the three-fold balanced product tricycle code can be equipped with a natural cup-product if the preorientiations of the classical codes obey certain conditions. These conditions are listed in proposition 5.2 of [16] which we repeat here in a form that is adapted for group-algebra codes. Below, we will use $|a|$ to measure the weight of the group-algebra element $a$ – i.e the number of group elements that appear in its expansion. We also write $a_{1}\cap a_{2}$ for $a_{1},a_{2}\in\mathbb{F}_{2}[G]$ to denote the set of group elements that are common to both group-algebra elements. The conditions on the classical preorientation of the code defined by $a\in\mathbb{F}_{2}[G]$ are then (with the conditions on the $b$ and $c$ classical code preorientations defined identically):
$|a_{in}|+|a_{out}|=0\ mod(2)$ (22)
$|a_{in}\cdot g\cap a_{free}\cdot h|=0\ mod(2)\quad\forall g,h\in G:\ g\neq h$ (23)
$|a_{out}\cdot g\cap a_{free}\cdot h|=0\ mod(2)\quad\forall g,h\in G:\ g\neq h$ (24)
$|a_{out}\cdot f\cap a_{free}\cdot g\cap a_{in}\cdot h|=0\ mod(2)\quad\forall f,g,h\in G:\ f\neq g\neq h,\ f\neq g$ (25)
where the multiplication in the expressions above is in the group-algebra sense. For three classical group-algebra codes with preorientiations that satisfy the above conditions, the balanced product yields a cup-product on the quantum code’s cochain complex, and consequently produces a constant depth $CCZ$ circuit of the form Eq.˜20. We will not reproduce the details of how the quantum code’s cup-product follows from the above pre-orientations, and refer the reader to [16] for details.
We call a preorientation non-trivial if either the ‘in’ or ‘out’ partitions are non-zero, since only these terms contribute to the cup-product – the ‘free’ partition has no effect besides excluding certain group elements from contributing. Our first technical result on cup-products for the cochain complexes of the $(w_{X},w_{Z})=(12,8)$ check-weight tricycle codes we consider in this work is that any valid, non-trivial preorientation of the classical codes must have the ‘free’ partition equal to 0. This follows from the next theorem.
@@ Appendix B Cup products and transversal CCZ action
A classical group-algebra code defined by a weight-44 group-algebra element $a\in\mathbb{F}_{2}[G]$ only has non-trivial preorientations satisfying the conditions in Eq.˜22-Eq.˜25 if $a_{free}=0$.
@@
Suppose $a_{free}\neq 0$ for the sake of contradiction. Since $|a|=4$, Eq.˜22 implies that the only non-trivial possibilities are $|a_{in}|=0,1,2,4$. If $|a_{in}|=4$ then the statement follows trivially. If $|a_{in}|=1$, Eq.˜22 implies that $|a_{out}|=1,|a_{free}|=2$. In this case, let $a_{in}=x$ and $a_{free}=y+z$, where $x,y,z\in G$. By assumption of disjointness of the partitions, $x\neq y\neq z$. Equation Eq.˜23 is now impossible to satisfy, as for any $g\in G$, $gx=hy$ for $h=gxy^{-1}$ and $gx=hz$ for $h=gxz^{-1}$. Thus, the RHS of Eq.˜23 is always $1$ if $a_{free}\neq 0$. Similar reasoning holds if $|a_{in}|=2$. If $a_{in}=0$, the same logic applies to $a_{out}$ using Eq.˜24 instead. ∎
A consequence of this theorem is that for the tricycle codes considered in this paper, we may restrict our attention to preorientations of the classical codes with empty free partitions. The condition in Eq.˜22 then implies that the only possibilities are $|a_{in}|=1$, $|a_{out}|=3$ (or vice-versa) and $|a_{in}|=|a_{out}|=2$. Any such partition of $a$ thus produces a valid non-trivial preorientiation, as the remaining conditions Eq.˜23- Eq.˜25 are trivially satisfied by empty ‘free’ partitions.
Next, we define the $f_{CCZ}$ function for the quantum tricycle codes that arises from the cup product specified by a choice of preorientations on the constituent classical group-algebra codes. The structure of the cochain complex in Eq.˜18 implies that we can partition the qubits of the code into three sectors of $|G|$ qubits each – see Appendix˜A. Within each sector, a qubit can be labelled by a group element. Given some order of the group elements, the notation $g^{i}$ will thus be used to denote qubit $g$ of the $i$-th sector ($i=I,II,III$). For the following, denote $\alpha_{I}\equiv a$, $\alpha_{II}\equiv b$, $\alpha_{III}\equiv c$ as the three group-algebra elements that specify the classical codes (this notation makes the expression below compact).
Proposition 2.
$f_{CCZ}(p^{i},q^{j},r^{k})=|p^{-1}\alpha_{i}^{in}\cap q^{-1}\alpha_{j}^{out}|\cdot|q^{-1}\alpha_{j}^{in}\cap r^{-1}\alpha_{k}^{out}|\cdot\delta_{i\neq j\neq k}\quad\mathrm{mod}(2)$ (26)
Proof (informal).
This definition follows from a straightforward but tedious computation of the cup-product on the balanced-product quantum code induced by the classical code cup-products. See section 5.3.2 of [16] for the case $i=I,j=II,k=III$. The other $5$ possible orderings of $i,j,k$ follow a similar computation, which can be combined into the compact expression in Eq.˜26. ∎
Note that the expression in Proposition˜2 holds for any tricycle code, not just those constructed from weight-$4$ group-algebra elements. The $\delta_{i\neq j\neq k}$ structure is manifest in the illustrations of the circuits in Fig.˜1, where no $CCZ$ acts between qubits in the same sector across code-blocks. We can now elucidate how the ‘single-sector’ (SS) and ‘triple-sector’ (TS) circuits introduced in Sec.˜2.2 emerge from the structure of $f_{CCZ}$. A single-sector circuit acting across sector $I$ of block 1, sector $II$ of block 2, and sector $III$ as in Fig.˜1 b) is constructed by choosing $\alpha_{I}^{in}=\alpha_{I}$ and $\alpha_{III}^{out}=\alpha_{III}$, with other partitions of $\alpha_{I}$ and $\alpha_{III}$ set to 0 and both partitions of $\alpha_{II}$ non-zero. This makes $f_{CCZ}=0$ on any other combination of argument sectors. A triple-sector circuit instead has all ‘in’ and ‘out’ partitions of $\alpha_{I},\alpha_{II},\alpha_{III}$ non-zero. In a triple-sector circuit, gates involved between sectors IIIIIII-II-III, IIIIIIII-III-I, and IIIIIIIII-I-II of the three blocks can be parallelized, while sectors IIIIIII-III-II, IIIIIIII-I-III, and IIIIIIIII-II-I can be done in another set of parallelizable gates – these correspond to the orange and black gate layers in Fig.˜1 c).
Using the expression in Eq.˜26, we can analyze the depth of the resulting CCZ circuits.
Theorem B.2.
All tricycle codes constructed from weight4-4 group-algebra elements $\alpha_{I}$, $\alpha_{II}$, $\alpha_{III}$ have depth-$18$ code-space preserving $CCZ$ circuits.

### v2

@@ Appendix A Code construction and properties
This homological viewpoint is especially useful for constructing transversal logical gates. In the next section, we apply the formalism of [18] to equip the complex Eq.˜25 with an algebraic structure that enables transversal CCZCCZ circuits.
Appendix B Single-shot guarantees
We prove that tricycle codes possess soundness properties in one basis, which in turn guarantee some fault-tolerant single-shot state preparation and error correction capabilities under adversarial noise. Our proofs heavily reference Campbell’s results on the soundness of 4D homological product codes in Ref. [22].
Definition 4.
(Soundness): Let $t\in\mathbb{Z}_{+}$ and $f:\mathbb{Z}\rightarrow\mathbb{R}$ be a function with $f(0)=0$. For some ring $R$ and norm $|\cdot|$ on $R^{n}$, we say that a parity check matrix $H\in R^{m\times n}$ is a $(t,f)$-sound map if, for any $v\in R^{n}$ such that $|Hv|<t$, there exists some $v^{\prime}\in R^{n}$ such that $Hv=Hv^{\prime}$ and $|v|^{\prime}\leq f(|Hv|)$.
To begin, we consider homological product codes formed from classical seed codes over $\mathbb{F}_{2^{k}}$ for some $k>1$. Let $|\cdot|$ denote the Hamming weight: for $v\in\mathbb{F}_{2^{k}}^{n}$, $|v|=|\{i:v_{i}\neq 0\}|$. A similar Hamming weight applies over $R$. For a classical code over $\mathbb{F}_{2^{k}}$-vector spaces, we define its distance in the usual way with respect to the Hamming weight.
Lemma B.1.
(Extension to the first soundness lemma, [22]): Let $\mathcal{C}=C^{0}\xrightarrow{\delta^{0}}C^{1}$ be a cochain complex over $\mathbb{F}_{2^{k}}$-vector spaces for some $k>1$ corresponding to a code over $l$ dits. Denote the Hamming distance of $\mathcal{C}$ by $d_{H}(\delta^{0})=d_{0}$ and the transpose code by $d_{0}^{T}$. Let $\tilde{\mathcal{C}}=\mathcal{C}\otimes_{\mathbb{F}_{2^{k}}}\mathcal{C}$, yielding the $3$-term complex $\tilde{\mathcal{C}}=\tilde{C}_{-1}\xrightarrow{\tilde{\delta}^{-1}}\tilde{C}^{0}\xrightarrow{\tilde{\delta}^{0}}\tilde{C}^{1}$, where
$\tilde{\delta}^{0T}=\begin{pmatrix}\delta^{0T}\otimes_{\mathbb{F}_{2^{k}}}I\\ I\otimes_{\mathbb{F}_{2^{k}}}\delta^{0}\end{pmatrix},\qquad\ \tilde{\delta}^{-1}=\begin{pmatrix}I\otimes_{\mathbb{F}_{2^{k}}}\delta^{0T}\\ \delta^{0}\otimes_{\mathbb{F}_{2^{k}}}I\end{pmatrix}$. (26)
Then, $\tilde{\delta}^{0T},\ \tilde{\delta}^{-1}$ are $(t,f)$-sound, where $t=\min\{d_{0},d_{0}^{T}\}$ and $f(x)=x^{2}/4$.
We provide a self-contained proof of this lemma, which mostly mirrors Appendix D in Ref. [22], for illustration. In later adaptations, we only discuss necessary modifications to results from Ref. [22].
Proof.
We focus on $\tilde{\delta}^{0T}$; the proof sketch for $\tilde{\delta}^{-1}$ is similar. Let $s\in\operatorname{\text{im}}\tilde{\delta}^{0T}$ with $|s|<t$. $v\in\mathbb{F}_{2^{k}}^{n}$, $s=\tilde{\delta}^{0T}v$. Letting $s=(s_{L}\ s_{R})$ and reshaping $\mathbb{F}_{2^{k}}^{l}\otimes_{\mathbb{F}_{2^{k}}}\mathbb{F}_{2^{k}}^{l}\rightarrow\mathbb{F}_{2^{k}}^{l\times l}$ yields the equations $S_{L}=\delta^{0T}V$ and $S_{R}=V\delta^{0T}$. For any $A\in\mathbb{F}_{2^{k}}^{l\times l}$, let $\textnormal{rsp}(A)\subseteq\{1,\dots,n\}$ denote the set of indices corresponding to nonzero rows of $A$, and likewise with $\textnormal{csp}(A)$ for columns.
Lemma B.2.
There exists some $w\in\mathbb{F}_{2^{k}}^{l}$ such that $s=\tilde{\delta}^{0T}w$, $|\textnormal{csp}(S_{L})|=|\textnormal{csp}(W)|$, and $|\textnormal{rsp}(S_{R})|=|\textnormal{csp}(W)|$.
Proof.
At least one $w$ satisfying the syndrome equation must exist. For any such $w$, $\textnormal{csp}(S_{L})\subseteq\textnormal{csp}(W)$ and $\textnormal{rsp}(S_{R})\subseteq\textnormal{rsp}(W)$, $|\textnormal{csp}(S_{L})|\leq|\textnormal{csp}(W)|$ and $|\textnormal{rsp}(S_{R})|\leq|\textnormal{rsp}(W)|$. The $\geq$ direction is more difficult to prove. For concreteness, assume $|\textnormal{csp}(S_{L})|<|\textnormal{csp}(V)|$ and $|\textnormal{rsp}(S_{R})|<|\textnormal{rsp}(V)|$. Then, there exist column and row vectors $a$ and $b^{T}$ in $V$ such that $a\in\ker\delta^{0T}$ and $b\in\ker\delta^{0}$. $b^{T}$ is the $i$th row of $V$, which we denote by $V_{i,\cdot}=b^{T}$, and $a$ is the $j$th column: $V_{\cdot,j}=a$. If $V_{ij}\neq 0$, then $V^{\prime}=V+V_{ij}^{-1}ab^{T}$ has strictly smaller support size: $V^{\prime}_{\cdot,j}=V_{\cdot,j}+V_{ij}^{-1}(ab^{T})_{\cdot,j}=0$ since $(ab^{T})_{\cdot,j}=b^{T}_{j}a=V_{ij}a$. Identically, $V^{\prime}_{i,\cdot}=0$, and no new rows or columns can pick up support. Thus, we update $V\leftarrow V^{\prime}$ and repeat until the algorithm stalls, which is when there are no rows and columns $b^{T}$ and $a$ such that $V_{ij}\neq 0$. Up to row and column permutations, this leaves $V$ in the form
$V=\begin{pmatrix}D&B&0\\ A&C&0\\ 0&0&0\end{pmatrix}$ (27)
where columns in the left column block contain $a\in\ker\delta^{0T}$, and those in the middle column block satisfy $\delta^{0T}a\neq 0$. Similarly, rows $b^{T}$ in the top row block satisfy $b\in\ker\delta^{0}$, whereas for those in the middle block, $\delta^{0}b\neq 0$. Since no row-column pairs intersect, $D=0$. If $A$ is nonzero, any nonzero column $a\in\ker\delta^{0T}$ satisfies $|a|\geq d_{0}^{T}$, $|\text{rsp}(S_{R})|\geq|c|\geq d_{0}^{T}$, which is a contradiction. Thus, $A=0$, and similarly, $B=0$. ∎
Since $|s|=|S_{L}|+|S_{R}|\geq|\text{csp}(S_{L})|+|\text{rsp}(S_{R})|$, the general identity $(a+b)^{2}/4\geq ab$ for integers $a,b$ yields
$|s|^{2}/4\geq|\text{csp}(S_{L})||\text{rsp}(S_{R})|=|\text{csp}(V)||\text{rsp}(V)|\geq|V|.$ (28)
Similarly, we have the following result.
Lemma B.3.
(Extension to the second soundness lemma, [22]): Let $\tilde{C}^{-1}\xrightarrow{\tilde{\delta}^{-1}}\tilde{C}^{0}\xrightarrow{\tilde{\delta}^{0}}\tilde{C}^{1}$ be a cochain complex over $\mathbb{F}_{2^{k_{i}}}$-vector spaces where $\tilde{\delta}^{0T}$, $\tilde{\delta}^{-1}$ are $(t,f)$-sound with $f(x)=x^{2}/4$. Then, $\breve{\mathcal{C}}=\tilde{\mathcal{C}}\otimes_{\mathbb{F}_{2^{k}}}\tilde{\mathcal{C}}$ is the 4-complex
$\breve{\mathcal{C}}=\breve{C}^{-2}\xrightarrow{\breve{\delta}^{-2}}\breve{C}^{-1}\xrightarrow{\breve{\delta}^{-1}}\breve{C}^{0}\xrightarrow{\breve{\delta}^{0}}\breve{C}^{1}\xrightarrow{\breve{\delta}^{1}}\breve{C}^{2}$ (29)
where
$\check{\delta}^{-2}=\begin{pmatrix}I\otimes_{\mathbb{F}_{2^{k}}}\tilde{\delta}_{0}^{T}\\ \tilde{\delta}_{-1}\otimes_{\mathbb{F}_{2^{k}}}I\end{pmatrix},\qquad\check{\delta}^{-1}=\begin{pmatrix}I\otimes_{\mathbb{F}_{2^{k}}}\tilde{\delta}_{-1}^{T}&0\\ \tilde{\delta}_{-1}\otimes_{\mathbb{F}_{2^{k}}}I&I\otimes_{\mathbb{F}_{2^{k}}}\tilde{\delta}_{0}^{T}\\ 0&\tilde{\delta}_{0}\otimes_{\mathbb{F}_{2^{k}}}I\end{pmatrix}$, (30)
$\check{\delta}^{0T}=\begin{pmatrix}\tilde{\delta}_{-1}^{T}\otimes_{\mathbb{F}_{2^{k}}}I&0\\ I\otimes_{\mathbb{F}_{2^{k}}}\tilde{\delta}_{-1}&\tilde{\delta}_{0}^{T}\otimes_{\mathbb{F}_{2^{k}}}I\\ 0&I\otimes_{\mathbb{F}_{2^{k}}}\tilde{\delta}_{0}\end{pmatrix},\qquad\check{\delta}^{1T}=\begin{pmatrix}\tilde{\delta}_{0}^{T}\otimes_{\mathbb{F}_{2^{k}}}I\\ I\otimes_{\mathbb{F}_{2^{k}}}\tilde{\delta}_{-1}\end{pmatrix}$. (31)
The maps $\breve{\delta}^{0}$ and $(\breve{\delta}^{-1})^{T}$ are $(t,g)$-sound, with $g(x)=x^{3}/4$.
Proof.
For $k=1$, the proof of this result is Lemma 6 in [22], which in turn relies on “partial soundness” and “inheritance of soundness” results, which are Lemma 7 and Claim 3 there, respectively. For $k>2$, the proofs of all three statements can be adapted with a small modification, just like lemma˜B.2: during the reduction steps $V\leftarrow V+ab^{T}$ is modified into $V\leftarrow V+V_{ij}^{-1}ab^{T}$. The $\mathbb{F}_{2}$ case is simply the special case where $V_{ij}^{-1}=1$ given $a$ and $b^{T}$ intersect. Still working within the Hamming norm, the remainder of the analysis extends. ∎
We can invoke these to demonstrate soundness properties for some connected 4D balanced product codes.
@@ Appendix B Single-shot guarantees
Let $\breve{\mathcal{C}}$ denote a 4-term cochain complex Eq.˜29 over $R$-modules, built from the classical seed code $\mathcal{C}=C^{0}\xrightarrow{\delta^{0}}C^{1}$ by repeated applications of the $R$-valued homological product, where $R=\mathbb{F}_{2}[N]$. Assume $|N|$ is odd. Let $d_{0}$ denote the distance of the binarized classical code $\mathbbm{b}_{N}(C^{0})\xrightarrow{\mathbb{B}_{N}(\delta^{0})}\mathbbm{b}_{N}(C^{1})$. Then, $\breve{C}$ has $(t,f)$-soundness, where $t=d_{0}/|N|$, and $g(x)=\frac{|N|+1}{8}x^{3}$
@@
Since $|N|$ is odd, $|N|\nmid\text{char}(\mathbb{F}_{2})=2$, so by Maschke’s theorem, $R$ is semisimple. From the Wedderburn-Artin theorem, it can be shown that $R\cong\mathbb{F}_{2^{k_{1}}}\times\cdots\times\mathbb{F}_{2^{k_{m}}}$ [77]. The explicit form of the isomorphism is $\psi:r\mapsto(r\cdot e_{1},r\cdot e_{2},\dots,r\cdot e_{m})$ where $\{e_{1},\dots,e_{m}\}$ is the set of primitive orthogonal idempotents, which also satisfy a completeness relation $\sum_{i}e_{i}=1_{R}=1_{G}$. The isomorphism can be extended to matrix rings $R^{m\times n}\cong\bigoplus_{i}\mathbb{F}_{2^{k_{i}}}^{m\times n}$. Let $A\in R^{n\times m}$ and $B\in R^{m\times p}$. Then, $AB=\bigoplus_{i,j}Ae_{i}Be_{j}=\bigoplus_{i}A_{(k_{i})}B_{(k_{i})}$, and, identically, $A\otimes_{R}B=\bigoplus_{i}[A_{(k_{i})}\otimes_{\mathbb{F}_{2^{k_{i}}}}B_{(k_{i})}]$. Therefore, $\mathcal{C}\simeq\bigoplus_{i}\mathcal{C}_{(k_{i})}$, where $\mathcal{C}_{k_{i}}$ is a $\mathbb{F}_{2^{k_{i}}}$-code, with the parity check matrix decomposed as $\delta^{0}\simeq\bigoplus_{i}\delta^{0}_{(k_{i})}$.
Note that for any $v\in R^{m}$, $|v|_{R}\leq|N||v|$. Thus, $d_{H}(\delta^{0}_{(k_{i})})\geq d_{H}(\delta^{0})\geq d_{R}(\delta^{0})/|N|=d_{H}(\mathbb{B}_{N}(\delta_{0}))/|N|$ provides a lower bound on the Hamming distance of each $\delta^{0}_{(k_{i})}$. Since the tensor product over $R$ splits component-wise, one can see the $R$-homological product codes also decompose component-wise, e.g., $\breve{\mathcal{C}}=\bigoplus_{i}[\tilde{\mathcal{C}}_{(k_{i})}\otimes_{\mathbb{F}_{2^{k_{i}}}}\tilde{\mathcal{C}}_{(k_{i})}]$. We apply lemmas B.1 and B.3 to each classical code component $\mathcal{C}_{(k_{i})}$, obtaining that every $\breve{\delta}^{0}_{(k_{i})}$ (and $(\breve{\delta}^{-1}_{(k_{i})})^{T}$) are at least $(d_{H}(\delta^{0}_{(k_{i})}),g)$-sound. Let $s\in\operatorname{\text{im}}\tilde{\delta}_{0}$ decompose as $s=\bigoplus_{i}s_{(k_{i})}$. Since $\psi$ is any isomorphism, so $\ker\psi^{-1}=\{0\}$, $|s_{(k_{i})}|_{H}\leq|s|_{H}$. If $|s|_{R}\leq d_{H}(\mathbb{B}_{N}(\delta_{0}))/|N|$, then it follows by soundness that for each $i$, there exists $v_{(k_{i})}\in\mathbb{F}_{2^{k_{i}}}^{6l^{4}}$ such that $s_{(k_{i})}=\breve{\delta}^{0}_{(k_{i})}v_{(k_{i})}$, and $|v_{(k_{i})}|\leq|s_{(k_{i})}|^{3}/4$. This gives $|v|_{R}\leq m|v_{(k_{i})}|_{H}\leq m|s|_{H}^{3}/4\leq m|s|^{3}_{R}/4=m|\mathbbm{b}_{N}(s)|^{3}/4$. Finally, since $m$ is the number of equivalence classes of integer pairs $(x,y)$ such that $y=2^{k}x\bmod|N|$ for some integer $k$ (formally known as the number of 2-cyclotomic cosets modulo $|N|$), it has the upper bound $m\leq(|N|+1)/2$. ∎
Corollary B.4.1.
(Soundness of tricycle codes): Consider any connected tricycle code with parameters $n,k,(d_{Z},d_{X})]]$, which can be expressed as the balanced product of binary classical codes $\mathcal{C}_{1},\mathcal{C}_{2}$, and $\mathcal{C}_{3}$. There exists a decoder such that for any pattern of measurement errors $u\in\mathbb{F}_{2}^{n}$ satisfying $|u|\leq\frac{1}{2|N|}\min_{i}d_{i}$ and data errors $v\in\mathbb{F}_{2}^{n}$ satisfying $(|N|+1)|u|+|v|<d_{Z}/2$, the decoder takes in the syndrome $H_{Z}v+u$ and outputs $\hat{v}$ such that, modulo stabilizers, $|(\hat{v}+v)\bmod 2|$ leaves a residual error of size at most $(|N|+1)|u|$.